# V2V-LLM: Vehicle-to-Vehicle Cooperative Autonomous Driving with Multi-Modal Large Language Models

### v3

@@ Abstract
Current autonomous driving vehicles rely mainly on their individual sensors to understand surrounding scenes and plan for future trajectories, which can be unreliable when the sensors are malfunctioning or occluded. To address this problem, cooperative perception methods via vehicle-to-vehicle (V2V) communication have been proposed, but they have tended to focus on perception tasks like detection or tracking. How those approaches contribute to overall cooperative planning performance is still under-explored. Inspired by recent progress using Large Language Models (LLMs) to build autonomous driving systems, we propose a novel problem setting that integrates a Multi-Modal LLM into cooperative autonomous driving, with the proposed Vehicle-to-Vehicle Question-Answering (V2V-QA) dataset and benchmark. We also propose our baseline method Vehicle-to-Vehicle Multi-Modal Large Language Model (V2V-LLM), which uses an LLM to fuse perception information from multiple connected autonomous vehicles (CAVs) and answer various types of driving-related questions: grounding, notable object identification, and planning. Experimental results show that our proposed V2V-LLM can be a promising unified model architecture for performing various tasks in cooperative autonomous driving, and outperforms other baseline methods that use different fusion approaches. Our work also creates a new research direction that can improve the safety of future autonomous driving systems. The code and data will be released to the public to facilitate open-source research in this field.
@@
1 Introduction
Autonomous driving technology has advanced significantly due to the evolution of deep learning algorithms, computing infrastructures, and the release of large-scale real-world driving datasets and benchmarks [13, 3, 38]. However, the perception and planning systems of autonomous vehicles in daily operation rely mainly on their local LiDAR sensors and cameras to detect notable nearby objects and plan for future trajectories. This approach may encounter safety-critical problems when the sensors are occluded by nearby large objects. In such situations, autonomous driving vehicles are unable to accurately detect all nearby notable objects, making subsequent planning results unreliable.
Other recent research has attempted to use LLM-based methods to build end-to-end perception and planning algorithms for an individual autonomous vehicle [37, 40, 43, 39, 5, 34, 46, 54] due to their common-sense reasoning and generalization ability from large-scale pre-trained data. These LLM-based models encode the raw sensor inputs into visual features, and then perform visual understanding and answer varieties of driving-related perception and planning questions. These approaches have shown some promise but have not yet explored the benefits of cooperative perception and planning. LLM-based driving algorithms without cooperative perception could also face safety-critical issues when the individual vehicle’s sensory capability is limited.
In this paper, we propose and explore a novel problem setting wherein LLM-based methods are used to build end-to-end perception and planning algorithms for Cooperative Autonomous Driving, as illustrated in Fig. 1. In this problem setting, we assume that there are multiple CAVs and a centralized LLM computing node. All CAVs share their individual perception information with the LLM. Any CAV can ask the LLM a question in natural language to obtain useful information for driving safety. To enable the study of this problem setting, we first create the Vehicle-to-Vehicle Question-Answering (V2V-QA) dataset, built upon the V2V4Real [53] and V2X-Real [48] cooperative perception datasets for autonomous driving. Our V2V-QA  includes grounding (Figs. 2(a), 2(b) and 2(c)), notable object identification (Fig. 2(d)), and planning (Fig. 2(e)) question-answer pairs. There are several differences between our novel problem setting and other existing LLM-based driving research [39, 43, 37, 33, 35, 4]. First, our LLM can fuse multiple perception information from different CAVs and provide answers to different questions from any CAV, rather than just serving a single self-driving car. Second, our grounding questions are specially designed to focus on the potential occluded regions of each individual CAV. More differences between our V2V-QA and other related datasets are summarized in Tab. 1.
Dataset Publication # CAVs Sim/Real # Frames # QA # QA/frame Point Cloud Planning
AD
NuScenes [3] CVPR 2020 - Real 400K - -
Waymo [38] CVPR 2020 - Real 200K - -
@@ 1 Introduction
V2V4Real [53] CVPR 2023 2 Real 20K - -
@@
NuScenes-QA [35] AAAI 2024 - Real 34K 460K 13.5
Lingo-QA [33] ECCV 2024 - Real 28K 420K 15.3
MAPLM-QA [4] CVPR 2024 - Real 14K 61K 4.4
DriveLM [37] ECCV 2024 - Sim+Real 69K 2M 29.1
TOKEN [39] CoRL 2024 - Real 28K 434K 15.5
OmniDrive-nuScenes [43] arXiv 2024 - Real 34K 450K 13.2
@@
Table 1: Comparison between our V2V-QA and recent related Autonomous Driving (AD) datasets.
To establish a benchmark for the V2V-QA dataset, we first propose a strong baseline method: Vehicle-to-Vehicle Multi-Modal Large Language Model (V2V-LLM) for cooperative autonomous driving, as illustrated in Fig. 3. Each CAV extracts its own perception features and shares them with V2V-LLM. The V2V-LLM fuses the scene-level feature maps and object-level feature vectors, and then performs vision and language understanding to provide the answer to the input driving-related questions in V2V-QA. We also compare V2V-LLM with other baseline methods corresponding to different feature fusion methods: no fusion, early fusion, and intermediate fusion [53, 48, 52, 51, 50]. The results show that V2V-LLM achieves the best performance in the more important notable object identification and planning tasks and competitive performance in the grounding tasks, achieving strong performance for the overall autonomous driving system.
@@
  • We propose a baseline method V2V-LLM for cooperative autonomous driving to provide an initial benchmark for V2V-QA. This method fuses scene-level feature maps and object-level feature vectors provided by multiple CAVs, and answers different CAV’s driving-related questions.
@@
  • We create a benchmark for V2V-QA and show that V2V-LLM outperforms other baseline fusion methods on the notable object identification and planning tasks and achieves competitive results on the grounding tasks, indicating the potential of V2V-LLM to be a foundation model for cooperative autonomous driving.
2 Related Work
2.1 Cooperative Perception in Autonomous Driving
@@
Cooperative perception [16] algorithms were proposed to address the potential occlusion problem in individual autonomous vehicles. Pioneering work F-Cooper [6] proposes the first intermediate fusion approach that merges feature maps to achieve good cooperative detection performance. V2VNet [44] builds graph neural networks for cooperative perception. DiscoNet [23] adopts a knowledge distillation approach. More recent work, AttFuse [52], V2X-ViT [51], and CoBEVT [50] integrate attention-based models to aggregate features. Another group of works [29, 15, 55, 8] focuses on developing efficient communication approaches.
@@
From a dataset perspective [28, 57], simulation datasets, OPV2V [52], V2X-Sim [24], and V2XSet [51] were first generated for cooperative perception research. More recently, real datasets have been collected. V2V4Real [53] is the first worldwide available real vehicle-to-vehicle cooperative perception dataset with perception benchmarks. V2X-Real [48], DAIR-V2X [59], and TUMTraf-V2X [60] further include sensor data from roadside infrastructures.
@@
Different from this group of research, our problem setting and proposed V2V-QA dataset include both perception and planning question-answering tasks for multiple CAVs. Our proposed V2V-LLM also adopts a novel LLM-based fusion approach.
2.2 LLM-based Autonomous Driving
@@
Language-based planning models [31, 32, 21] first transforms the driving scene, object detection results, and ego-vehicle’s state into text input to the LLM. Then the LLM generates text output including the suggested driving action or the planned future trajectory. More recent approaches [39, 43, 37, 40, 54, 45, 56] use Multi-Modal Large Language Models (MLLMs) [25, 26, 27, 22, 17, 36, 41, 42, 1, 2] to encode point clouds or images into visual features. Then, the visual features are projected to the language embedding space for LLM to perform visual understanding and question-answering tasks.
@@
From a dataset perspective, several LLM-based autonomous driving datasets have been built on top of existing autonomous driving datasets. For example, Talk2Car [11], NuPrompt [47], NuScenes-QA [35], NuInstruct [49], and Reason2Drive [34] create captioning, perception, prediction, and planning QA pairs based on the NuScenes [3] dataset. BDD-X [18] is extended from BDD100K [58]. DriveLM [37] adopts real data from NuScenes [3] and simulated data from CARLA [12] to have larger-scale and more diverse driving QAs. Other datasets curated independently focus on different types of QA tasks. HAD [19] contains human-to-vehicle advice data. DRAMA [30] introduces joint risk localization and captioning. Lingo-QA [33] proposed counterfactual question-answering tasks. MAPLM-QA [4] emphasizes map and traffic scene understanding.
@@
Different from those LLM-based driving research that only supports individual autonomous vehicles, our problem setting and proposed V2V-QA dataset are designed for cooperative driving scenarios with multiple CAVs. In our problem setting, the LLM can aggregate perception features from multiple CAVs and provide answers to questions from different CAVs. Our V2V-QA is also designed to focus on the potential occluded regions. In addition, our V2V-QA contains both highway and urban cooperative driving scenarios, making our planning task more challenging than prior works based on the NuScenes [3] dataset.
3 V2V-QA Dataset
@@
3.1 Problem Setting
Our proposed V2V cooperative autonomous driving with LLM problem is illustrated in Fig. 1. In this setting, we assume there are multiple Connected Autonomous Vehicles (CAVs) and a centralized LLM computing node. All CAVs share their individual perception information, such as scene-level or object-level features, with the centralized LLM. Any CAV can ask the LLM a question in natural language to obtain information for driving safety. The LLM aggregates the received perception information from multiple CAVs and provides a natural language answer to the CAV’s question. In this research, the questions and answers include grounding (Q1-3), notable object identification (Q4), and planning (Q5), as illustrated in Fig. 2.
3.2 Dataset Details
Our V2V-QA dataset contains two splits: V2V-split and V2X-split, which are built on top of V2V4Real [53] and V2X-Real [48] datasets, respectively. These base datasets are collected by driving two vehicles with LiDAR sensors simultaneously near to each other. These datasets also includes 3D bounding box annotations for other objects in the driving scenes. In V2V4Real [53], the training set has $32$ driving sequences and a total of $7105$ frames of data per CAV, and the testing set has $9$ driving sequences and a total of $1993$ frames of data per CAV. In V2X-Real [48], the training set has $43$ driving sequences and a total of $5772$ frames of data per CAV, and the testing set has $9$ driving sequences and a total of $1253$ frames of data per CAV. The frame rate is $10$Hz. In V2X-Real [48], some driving scenes also provide lidar point clouds from roadside infrastructures. In our V2X-split, we also include them as perception inputs to the LLM with the same approach as using CAVs’ lidar point clouds to answer CAVs’ questions. We follow the same training and testing settings from V2V4Real [53] and V2X-Real [48] when building our V2V-split and V2X-split.
Tab. 2 summarizes the numbers of QA pairs in our proposed V2V-QA’s V2V-split and V2X-split. We have $1.45$M QA pairs in total and $30.2$ QA pairs per frame on average. More details can be found in the supplementary materials.
@@ 3.2 Dataset Details
Table 2: Dataset statistics of our V2V-QA’s V2V-split and V2X-split. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. Q4: Notable object identification. Q5: Planning.
3.3 Question and Answer Pairs Curation
@@
For each frame of V2V4Real [53] and V2X-Real [48] datasets, we create $5$ different types of QA pairs, including $3$ types of grounding questions, $1$ type of notable object identification question, and $1$ type of planning question. These QAs are designed for cooperative driving scenarios. To generate instances of these QA pairs, we use V2V4Real [53] and V2X-Real [48]’s ground-truth bounding box annotations, each CAV’s ground-truth trajectories, and individual detection results as the source information. Then we use different manually designed rules based on the geometric relationship among the aforementioned entities and text templates to generate our QA pairs. The text template can be seen in Figs. 5 and 6. The generation rule of each QA type is described as follows.
@@
Q1. Grounding at a reference location (Fig. 2(a)): In this type of question, we ask the LLM to identify whether an object that occupies a specific query 2D location exists. If so, the LLM is expected to provide the center location of the object. Otherwise, the LLM should indicate that there is nothing at the reference location. We use the center locations of ground-truth boxes and every CAV’s individual detection result boxes as the query locations in the questions. By doing so, we can focus more on evaluating each model’s cooperative grounding ability on the potential false positive and false negative detection results.
@@
Q2. Grounding behind a reference object at a location (Fig. 2(b)): When a CAV’s field of view is occluded by a nearby large detected object, this CAV may want to ask the centralized LLM to determine whether there exists any object behind that occluding large object given the fused perception information from all CAVs. If so, the LLM is expected to return the object’s location and the asking CAV may need to drive more defensively or adjust its planning. Otherwise, the LLM should indicate that there is nothing behind the reference object. We use the center location of each detection result box as the query locations in these questions. We draw a sector region based on the relative pose of the asking CAV and the reference object, and select the closest ground-truth object in the region as the answer.
@@
Q3. Grounding behind a reference object in a direction (Fig. 2(c)): We further challenge the LLM on language and spatial understanding ability by replacing Q2’s reference 2D location with a reference directional keyword. We first get the closest detection result box in each of the $6$ directions of a CAV as the reference object. Then we follow the same approach in Q2 to get the closest ground-truth box in the corresponding sector region as the answer.
@@
Q4. Notable object identification (Fig. 2(d)): The aforementioned grounding tasks are intermediate tasks in the autonomous driving pipeline. More critical abilities of autonomous vehicles involve both identifying notable objects near planned future trajectories and adjusting future planning to avoid potential collisions. We extract $6$ waypoints from the ground-truth trajectory in the next $3$ seconds as the reference future waypoints in the questions. Then we get, at most, the $3$ closest ground-truth objects within $10$ meters of the reference future trajectory as the answer.
@@
Q5. Planning (Fig. 2(e)): Planning is important because the ultimate goal of autonomous vehicles is to navigate through complex environments safely and avoid any potential collision in the future. To generate the planning QAs, we extract $6$ future waypoints, evenly distributed in the next $3$ seconds, from each CAV’s ground-truth future trajectory as the answer. Our V2V-QA’s planning task is more challenging than other NuScenes [3]-based LLM-driving related works for a couple reasons. First, we support multiple CAVs in cooperative driving scenarios. The LLM model needs to provide different answers depending on which CAV is asking, while prior works only need to generate planning results for a single autonomous vehicle. Second, our V2V-QA’s ground-truth planning trajectories are more diverse. V2V-QA contains both urban and highway driving scenarios, while NuScenes [3] only includes urban driving scenarios. Detailed dataset statistics and comparison can be seen in the supplementary material.
3.4 Evaluation Metrics
We follow prior works [39, 43]’s approach to evaluate model performance. For the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), the evaluation metrics are F1 score, precision, and recall. The ground-truth answers and model outputs contain objects’ center locations. If the center distance between the ground-truth answer and the model output is less than a threshold value, this output is considered as a true positive. We set the threshold value to be $4$ meters, a typical length of a vehicle.
For the planning question (Q5), the evaluation metrics are L2 errors and collision rates. The ground-truth answers and model outputs contain $6$ future waypoints, so we calculate the L2 errors on those waypoints. When calculating collision rates, we assume each of the CAVs’ bounding box sizes is $4$ meters in length, $2$ meters in width, and $1.5$ meters in height. We place each CAV’s bounding box at the model’s output future waypoints and calculate the Intersection-over-Union (IOU) between the CAV’s bounding box and every ground-truth object’s annotation bounding box in those future frames. If the IOU is larger than 00, it is considered as a collision.

### v4

# V2V-LLM: Vehicle-to-Vehicle Cooperative Autonomous Driving with Multimodal Large Language Models

@@ Abstract
Current autonomous driving vehicles rely mainly on their individual sensors to understand surrounding scenes and plan for future trajectories, which can be unreliable when the sensors are malfunctioning or occluded. To address this problem, cooperative perception methods via vehicle-to-vehicle (V2V) communication have been proposed, but they have tended to focus on perception tasks like detection or tracking. How those approaches contribute to overall cooperative planning performance is still under-explored. Inspired by recent progress using Large Language Models (LLMs) to build autonomous driving systems, we propose a novel problem setting that integrates a Multimodal LLM into cooperative autonomous driving, with the proposed Vehicle-to-Vehicle Question-Answering (V2V-QA) dataset and benchmark. We also propose our baseline method Vehicle-to-Vehicle Multimodal Large Language Model (V2V-LLM), which uses an LLM to fuse perception information from multiple connected autonomous vehicles (CAVs) and answer various types of driving-related questions: grounding, notable object identification, and planning. Experimental results show that our proposed V2V-LLM  can be a promising unified model architecture for performing various tasks in cooperative autonomous driving, and outperforms other baseline methods that use different fusion approaches. Our work also creates a new research direction that can improve the safety of future autonomous driving systems. Our code and dataset are released to facilitate open-source research at https://eddyhkchiu.github.io/v2vllm.github.io/.
I Introduction
Autonomous driving technology has advanced significantly due to the evolution of deep learning algorithms, and the release of large-scale real-world driving datasets and benchmarks [9, 1, 24]. However, the perception and planning systems of autonomous vehicles in daily operation rely mainly on their local LiDAR sensors and cameras to detect notable nearby objects and plan for future trajectories. This approach may encounter safety-critical problems when the sensors are occluded by nearby large objects.
Other recent research has attempted to use LLM-based methods to build end-to-end perception and planning algorithms for an individual autonomous vehicle [23, 26, 27, 25, 2, 29] due to their common-sense reasoning and generalization ability from large-scale pre-trained data. These LLM-based models encode the raw sensor inputs and answer driving-related perception and planning questions. These approaches have shown some promise but have not yet explored the benefits of cooperative perception and planning.
@@ I Introduction
TABLE I: Comparison between our V2V-QA and recent related Autonomous Driving (AD) datasets.
Dataset Publication # CAVs RSU Sim/Real # Frames # QA # QA/frame Point Cloud Planning
@@
DAIR-V2X [38] CVPR 2022 1 Real 71K - -
V2V4Real [34] CVPR 2023 2 - Real 20K - -
TUMTrafV2X [39] CVPR 2024 1 Real 2K - -
@@
NuScenes-QA [21] AAAI 2024 - - Real 34K 460K 13.5
Lingo-QA [20] ECCV 2024 - - Real 28K 420K 15.3
DriveLM [23] ECCV 2024 - - Sim+Real 69K 2M 29.1
TOKEN [25] CoRL 2024 - - Real 28K 434K 15.5
OmniDrive [27] CVPR 2025 - - Real 34K 450K 13.2
@@
In this paper, we propose and explore a novel problem setting wherein LLM-based methods are used to build end-to-end perception and planning algorithms for Cooperative Autonomous Driving, as illustrated in Figure 1. In this problem setting, we assume that there are multiple CAVs and a centralized LLM computing node. All CAVs share their individual perception information with the LLM. Any CAV can ask the LLM a question in natural language to obtain useful information for driving safety. To enable the study of this problem setting, we first create the Vehicle-to-Vehicle Question-Answering (V2V-QA) dataset, built upon the V2V4Real [34] and V2X-Real [31] cooperative perception datasets for autonomous driving. Our V2V-QA  includes grounding , notable object identification , and planning question-answer pairs. There are several differences between our novel problem setting and other existing LLM-based driving research [25, 27, 23, 20, 21]. First, our LLM can fuse multiple perception information from different CAVs and provide answers to different questions from any CAV, rather than just serving a single self-driving car. Second, our grounding questions are specially designed to focus on the potential occluded regions of each individual CAV. More differences between our V2V-QA and other related datasets are summarized in Table I.
To establish a benchmark for the V2V-QA dataset, we first propose a strong baseline method: Vehicle-to-Vehicle Multimodal Large Language Model (V2V-LLM) for cooperative autonomous driving, as illustrated in Figure 3. Each CAV extracts its own perception features and shares them with V2V-LLM. The V2V-LLM fuses the scene-level feature maps and object-level feature vectors, and then performs vision and language understanding to provide the answer to the input driving-related questions in V2V-QA. We also compare V2V-LLM with other baselines corresponding to different feature fusion methods: no fusion, early fusion, and intermediate fusion [34, 31, 36, 35, 33]. The results show that V2V-LLM achieves better performance in the grounding, notable object identification, and planning tasks for the cooperative autonomous driving scenarios in general.
@@
  • We propose a baseline method V2V-LLM for cooperative autonomous driving to provide an initial benchmark for V2V-QA. This method fuses scene-level and object-level features provided by multiple CAVs, and answers different CAV’s driving-related questions.
@@
  • We create a benchmark for V2V-QA and show that V2V-LLM outperforms other baselines on the grounding, notable object identification, and planning tasks in general, indicating the potential of V2V-LLM to be a foundation model for cooperative autonomous driving.
II Related Work
II-A Cooperative Perception in Autonomous Driving
@@
Cooperative perception [11] algorithms were proposed to address the potential occlusion problem in individual autonomous vehicles. Pioneering work F-Cooper [3] proposes the first intermediate fusion approach that merges feature maps to achieve good cooperative detection performance. More recent work, AttFuse [36], V2X-ViT [35], and CoBEVT [33] integrate attention-based models to aggregate features for cooperative detection and tracking.
@@
From a dataset perspective [17, 37], simulation datasets OPV2V [36], V2X-Sim [14], and V2XSet [35] were first generated for cooperative perception research. More recently, real datasets have been collected. V2V4Real [34] is the first worldwide available real vehicle-to-vehicle cooperative perception dataset with perception benchmarks. V2X-Real [31], DAIR-V2X [38], and TUMTraf-V2X [39] further include sensor data from roadside infrastructures.
@@
Different from this group of research, our problem setting and proposed dataset include both perception and planning question-answering tasks for multiple CAVs. Our proposed V2V-LLM also adopts a novel LLM-based fusion approach.
II-B LLM-based Autonomous Driving
@@
Language-based planning models [18, 19, 13] and more recent Multimodal Large Language Model (MLLM)-based approaches [25, 27, 23, 26, 28] encode the driving scene and ego-vehicle’s state into text and visual features and use them as input to the LLM. Then the LLM generates text output including the suggested action or future trajectory.
@@
From a dataset perspective, several LLM-based autonomous driving datasets have been built on top of existing autonomous driving datasets. For example, NuPrompt [30], NuScenes-QA [21], and NuInstruct [32] create captioning, perception, prediction, and planning QA pairs based on the NuScenes [1] dataset. DriveLM [23] adopts real data from NuScenes [1] and simulated data from CARLA [8] to have larger-scale and more diverse driving QAs.
@@
Different from those LLM-based driving research that only supports individual autonomous vehicles, our problem setting and proposed V2V-QA dataset are designed for cooperative driving with multiple CAVs.
@@
III V2V-QA Dataset
III-A Problem Setting
Our proposed V2V cooperative autonomous driving with LLM problem is illustrated in Figure 1. In this setting, we assume there are multiple Connected Autonomous Vehicles (CAVs) and a centralized LLM computing node. All CAVs share their individual perception information, such as scene-level or object-level features, with the centralized LLM. Any CAV can ask the LLM a question in natural language to obtain information for driving safety. The LLM aggregates the received perception information from multiple CAVs and provides a natural language answer to the CAV’s question. In this research, the questions and answers include grounding (Q1-3), notable object identification (Q4), and planning (Q5), as illustrated in Figure 2.
III-B Dataset Details
Our V2V-QA dataset contains two splits: V2V-split and V2X-split, which are built on top of V2V4Real [34] and V2X-Real [31] datasets, respectively. These base datasets are collected by driving two vehicles with LiDAR sensors simultaneously near to each other. These datasets also includes 3D bounding box annotations for other objects in the driving scenes. In V2V4Real [34], the training set has $32$ driving sequences and a total of $7105$ frames of data per CAV, and the testing set has $9$ driving sequences and a total of $1993$ frames of data per CAV. In V2X-Real [31], the training set has $43$ driving sequences and a total of $5772$ frames of data per CAV, and the testing set has $9$ driving sequences and a total of $1253$ frames of data per CAV. The frame rate is 1010Hz. In V2X-Real [31], some driving scenes also provide LiDAR point clouds from roadside infrastructures. We also include them as perception inputs to the LLM with the same approach as using CAVs’ LiDAR point clouds. We follow the same training and testing settings from V2V4Real [34] and V2X-Real [31] when building our V2V-split and V2X-split. Table II summarizes the numbers of QA pairs in our proposed V2V-QA’s V2V-split and V2X-split. We have 1.451.45M QA pairs in total and $30.2$ QA pairs per frame on average.
TABLE II: Dataset statistics of our V2V-QA’s V2V-split and V2X-split. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. Q4: Notable object identification. Q5: Planning.
@@ III-B Dataset Details
III-C Question and Answer Pairs Curation
@@
For each frame of V2V4Real [34] and V2X-Real [31] datasets, we create $5$ different types of QA pairs, including $3$ types of grounding questions, $1$ type of notable object identification question, and $1$ type of planning question. These QAs are designed for cooperative driving scenarios. To generate instances of these QA pairs, we use V2V4Real [34] and V2X-Real [31]’s ground-truth bounding box annotations, each CAV’s ground-truth trajectories, and individual detection results as the source information. Then we use different manually designed rules based on the geometric relationship among the aforementioned entities and text templates to generate our QA pairs. The text template can be seen in Figures 4 and 5. The generation rule of each QA type is described as follows.
@@
Q1. Grounding at a reference location (2(a)): In this type of question, we ask the LLM to identify whether an object that occupies a specific query 2D location exists. If so, the LLM is expected to provide the center location of the object. Otherwise, the LLM should indicate that there is nothing at the reference location. We use the center locations of ground-truth boxes and every CAV’s individual detection result boxes as the query locations in the questions. By doing so, we can focus more on evaluating each model’s cooperative grounding ability on the potential false positive and false negative detection results.
@@
Q2. Grounding behind a reference object at a location (2(b)): When a CAV’s field of view is occluded by a nearby large detected object, this CAV may want to ask the centralized LLM to determine whether there exists any object behind that occluding large object given the fused perception information from all CAVs. If so, the LLM is expected to return the object’s location and the asking CAV may need to drive more defensively or adjust its planning. Otherwise, the LLM should indicate that there is nothing behind the reference object. We use the center location of each detection result box as the query locations in these questions. We draw a sector region based on the relative pose of the asking CAV and the reference object, and select the closest ground-truth object in the region as the answer.
@@
Q3. Grounding behind a reference object in a direction (2(c)): We further challenge the LLM on language and spatial understanding ability by replacing Q2’s reference 2D location with a reference directional keyword. We first get the closest detection result box in each of the $6$ directions of a CAV as the reference object. Then we follow the same approach in Q2 to get the closest ground-truth box in the corresponding sector region as the answer.
@@
Q4. Notable object identification (2(d)): The aforementioned grounding tasks are intermediate tasks in the autonomous driving pipeline. More critical abilities of autonomous vehicles involve both identifying notable objects near planned future trajectories and adjusting future planning to avoid potential collisions. We extract $6$ waypoints from the ground-truth trajectory in the next $3$ seconds as the reference future waypoints in the questions. Then we get, at most, the $3$ closest ground-truth objects within $10$ meters of the reference future trajectory as the answer.
@@
Q5. Planning (2(e)): Planning is important because the ultimate goal of autonomous vehicles is to navigate through complex environments safely and avoid any potential collision in the future. To generate the planning QAs, we extract $6$ future waypoints, evenly distributed in the next $3$ seconds, from each CAV’s ground-truth future trajectory as the answer. Our V2V-QA’s planning task is more challenging than other NuScenes [1]-based LLM-driving related works for a couple reasons. First, we support multiple CAVs in cooperative driving scenarios. The LLM model needs to provide different answers depending on which CAV is asking, while prior works only need to generate planning results for a single autonomous vehicle. Second, our V2V-QA’s ground-truth planning trajectories are more diverse. V2V-QA contains both urban and highway driving scenarios, while NuScenes [1] only includes urban driving scenarios.
III-D Evaluation Metrics
We follow prior works [25, 27]’s approach to evaluate model performance. For the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), the evaluation metric is F1 score. For the planning question (Q5), the evaluation metrics are L2 errors and collision rates.